OpenStudy (gschibby):

Another one for ya': 2xe^-x - x^2*e^-x = 0 Solve for x?

zepdrix (zepdrix):

\[\Large 2xe^{-x} - x^2e^{-x} = 0\]

OpenStudy (gschibby):

\[2xe^{-x}-x ^{2}e ^{-x}=0\]

zepdrix (zepdrix):

lol

OpenStudy (gschibby):

Thats the one yeah!

zepdrix (zepdrix):

Start by factoring out e^-x from each term:\[\Large e^{-x}(2x-x^2)=0\]Gives you something like that, yes?

OpenStudy (gschibby):

of course! Factoring, doh...

OpenStudy (gschibby):

But where do I go from there?

zepdrix (zepdrix):

Umm so recall the `Zero Factor Property`: If, \(\Large (a)(b)=0, \qquad a=0\qquad\text{and/or} \qquad b=0\) So we have 2 factors, we can set them each equal to zero to solve for x.\[\Large e^{-x}=0, \qquad\qquad 2x-x^2=0\]

zepdrix (zepdrix):

For that first one, you need to remember something important about exponentials! :O

OpenStudy (gschibby):

I'm a little rusty on those :P

zepdrix (zepdrix):